A Bank loaned out $12,000, part of it at the rate of 8% per year and the rest at the the rate of 18% per year. If the interest recieved totaled $1000, how much was loaned at 8%?
Answer by checkley71(8403)

.08X+.18(12000-X)=1000
.08X+2160-.18X=1000
-.1X=1000-2160
-.1X=-1160
X=-1160/-.1
X=11600 LOANED @ 8%
12000-11600=400 LOANED @ 18%
PROOF
.08*11600+.18*400=1000
928+72=1000
1000=1000